Please help me with this question:
Tickets at a basketball game were $4 for an adult and $2 for a student. The total number of tickets sold was 300. If the sales of tickets was $800, how many tickets of each were sold?
Thank you!

Found 2 solutions by Boreal, ikleyn:
Answer by Boreal(15235)

A+S=300
4A+2S=800
multiply top by -2 to eliminate S
-2A-2S=-600
4A+2S=800
2A=200
A=100 adult tickets @$4=$400
S=200 student tickets @$2=$400

Answer by ikleyn(52781)

.
Tickets at a basketball game were $4 for an adult and $2 for a student. The total number of tickets sold was 300.
If the sales of tickets was $800, how many tickets of each were sold?
Thank you!
~~~~~~~~~~~~~~~~~~~~

Two ways to solve.

Solution 1. Two equations

 A +  S = 300,   (1)
4A + 2S = 800.   (2)   (the "value" equation)

Divide eq.(2) by 2 (both sides). Rewrite you system as

 A + S = 300,    (1')
2A + S = 400.    (2')

Distract eq.(2') from (1') (both sides). You will get

A = 100.

Answer. 100 adults tickets and 300-100 = 200 student tickets.

Solution 2. One equation

Let "A" be the number of adults' tickets.
Then the number of students' tickets is 300-A.

The value equation is 

4A + 2*(300-A) = 800.

Simplify and solve for A.